OpenStudy (anonymous):

Find the length of the radius of a sphere with a surface area of 64.62 km. sq. A)1.125 B)2.25 C)2.27 D)3.375

OpenStudy (anonymous):

Use can use the formula : \[SA = 4\pi r^2\]

OpenStudy (anonymous):

SO HOW DO I SOLVE FOR R ?

OpenStudy (anonymous):

Or in terms of Radius : \[Radius = \sqrt{SA/4\pi}\]

OpenStudy (anonymous):

C ?

OpenStudy (anonymous):

Yes you did right..